Find the Jacobian ∂(x, y, z)/∂(u, v, w).

(a) u = xy, v = y, w = x + z (b) u = x + y + z, v = x + y − z, w = x − y + z


1
Expert's answer
2021-11-24T17:17:24-0500

Solution;

(a)

Given;

u=xy

v=y

w=x+z

Now;

"x=\frac uy=\frac uv"

"y=v"

"z=w-x=w-\frac uv"

"\frac{\partial(x,y,z)}{\partial(u,v,w)}=\begin{vmatrix}
 \frac{\partial x}{\partial u} &\frac{\partial x}{\partial v} &\frac{\partial x}{\partial w} \\
 \frac{\partial y}{\partial u} & \frac{\partial y}{\partial v}&\frac{\partial y}{\partial w}\\
\frac{\partial z}{\partial u}&\frac{\partial z}{\partial v}&\frac{\partial z}{\partial w}
\end{vmatrix}" ="\begin{vmatrix}
 \frac1v & \frac u{v^2}&0 \\
 0 & 1&0\\
-\frac1v&-\frac{u}{v^2}&1
\end{vmatrix}" ="\frac1v(1-0)-\frac{u}{v^2}(0-0)+0(0+\frac1v)"

Hence;

"\frac{\partial(x,y,z)}{\partial(u,v,w)}=\frac1v"

(b)

u=x+y+z

v=x+y-z

w=x-y+z

Now;

"v+w=x+y-z+x-y+z=2x"

Hence;

"x=\frac{v+w}{2}"

Also;

"u-v=x+y+z-x-y+z=2z"

Hence;

"z=\frac{u-v}{2}"

Also;

"y=x+z-w" ="\frac{v+w}{2}+\frac{u-v}{2}-w"

Hence;

"y=\frac{u-w}{2}"

Therefore;

"\frac{\partial(x,y,z)}{\partial(u,v,w)}=\begin{vmatrix}
 \frac{\partial x}{\partial u} &\frac{\partial x}{\partial v} &\frac{\partial x}{\partial w} \\
 \frac{\partial y}{\partial u} & \frac{\partial y}{\partial v}&\frac{\partial y}{\partial w}\\
\frac{\partial z}{\partial u}&\frac{\partial z}{\partial v}&\frac{\partial z}{\partial w}
\end{vmatrix}=\begin{vmatrix}
 0 &\frac 12&\frac12\\
 \frac12 & 0&-\frac{1}{2}\\
\frac12&-\frac12&0
\end{vmatrix}"="0(0+\frac14)-\frac12(0+\frac14)+\frac12(-\frac14-0)"

Therefore;

"\frac{\partial(x,y,z)}{\partial(u,v,w)}=-\frac14"